# GenUP: Generative User Profilers as In-Context Learners for Next POI Recommender Systems

### v1

@@ Abstract.
Traditional POI recommendation systems often lack transparency, interpretability, and scrutability due to their reliance on dense vector-based user embeddings. Furthermore, the cold-start problem—where systems have insufficient data for new users—limits their ability to generate accurate recommendations. Existing methods often address this by leveraging similar trajectories from other users, but this approach can be computationally expensive and increases the context length for LLM-based methods, making them difficult to scale. To address these limitations, we propose a method that generates natural language (NL) user profiles from large-scale, location-based social network (LBSN) check-ins, utilizing robust personality assessments and behavioral theories. These NL profiles capture user preferences, routines, and behaviors, improving POI prediction accuracy while offering enhanced transparency. By incorporating NL profiles as system prompts to LLMs, our approach reduces reliance on extensive historical data, while remaining flexible, easily updated, and computationally efficient. Our method is not only competitive with other LLM-based and complex agentic frameworks but is also more scalable for real-world scenarios and on-device POI recommendations. Results demonstrate that our approach consistently outperforms baseline methods, offering a more interpretable and resource-efficient solution for POI recommendation systems. Our source code is available at: https://github.com/w11wo/GenUP.
@@ 1. Introduction
Location-based social networks (LBSNs) generate vast amounts of data, offering significant potential for modeling human mobility patterns. One application of LBSN data is next Point-of-Interest (POI) recommendation, where the goal is to predict a user’s likely next destination based on their historical check-ins. POI recommendation systems have evolved from early Markov models (Cheng et al., 2013; Rendle et al., 2010) to more advanced deep learning techniques, such as sequential models (Kong and Wu, 2018; Sun et al., 2020) and Graph Neural Networks (GNNs) (Lim et al., 2020; Zhang et al., 2022; Rao et al., 2022; Yang et al., 2022a). These models have significantly improved the ability to capture spatial-temporal dependencies and user mobility patterns. However, challenges remain, particularly with the cold-start problem, where limited interaction data makes it difficult to provide accurate predictions for inactive users. This is typically resolved by learning from similar historical trajectories from other users. More recently, large language models (LLMs) have been applied to next POI recommendation (Wang et al., 2023; Beneduce et al., 2024), leveraging their commonsense reasoning capabilities and their ability to incorporate contextual information. However, LLM-based methods can be computationally intensive due to their reliance on long historical trajectory prompts to capture cross-user patterns (Li et al., 2024). Similarly, agentic LLM frameworks (Feng et al., 2024), which employ multiple agents for different tasks, introduce further complexity and resource demands, making them difficult to implement in real-time applications.
@@
Another challenge in POI recommendation is personalization, often achieved through dense vector representations of users. While these embeddings effectively model user behavior, they lack interpretability, transparency, and scrutability—qualities crucial for building user trust and improving recommendation explainability (O’Donovan and Smyth, 2005; Zhang and Chen, 2020). Natural language (NL) user profiles offer a promising alternative, where new users can quickly provide a concise description of their preferences, and existing users can update their profiles for more personalized and transparent recommendations (Radlinski et al., 2022). Despite their potential, NL profiles have not seen widespread adoption due to difficulties in evaluation and the ongoing development of methods for accurately generating these profiles (Ramos et al., 2024).
@@
In response to these limitations, we propose a method that introduces NL user profiles into next POI recommendation systems. By replacing long historical trajectory prompts with concise NL profiles, our approach significantly reduces computational overhead and shortens the context length required for LLM-based recommendations. Furthermore, it improves personalization by shifting from dense user vector representations to interpretable NL profiles. We propose generating these profiles by analyzing LBSN check-ins to learn user routines, preferences, and behaviors. To create more grounded NL profiles, we integrate robust behavioral theories, capturing underlying intentions and patterns for more accurate POI predictions. In summary, our contributions are as follows:
@@
  • We introduce natural language (NL) user profiles as system prompts to LLMs for next POI recommendation.
  • We leverage LBSN data to create NL user profiles, improving personalized recommendations and cross-user learning.
  • We propose a hybrid approach that combines the generation capabilities of LLMs with behavioral theories for more accurate user profiles.
@@ 2.1. Large Language Models for Next POI Recommendation
Large language models (LLMs) have since emerged as a viable solution for next POI recommendation, offering the ability to process both spatial-temporal data and commonsense knowledge. LLM-Mob (Wang et al., 2023), for example, introduced in-context learning and exploited the zero-shot reasoning capabilities of LLMs (Wei et al., 2022) to enhance next POI recommendation given the historical and contextual trajectories of a user. Similarly, LLM-ZS (Beneduce et al., 2024) simplified the zero-shot process and analyzed zero-shot, one-shot, and few-shot prompting scenarios. However, both methods depend heavily on task-specific prompting and require very large LMs, such as 70B instruction-tuned or closed-source LLMs, to achieve competitive performance. AgentMove (Feng et al., 2024) extends these techniques by employing multiple LLMs agents (Wang et al., 2024; Xi et al., 2023), each dedicated to different components of mobility prediction. This framework further integrates external knowledge sources such as urban structures, geospatial data, and location graphs, enabling the predictor LLM to process richer contextual information for the prediction process. While this multi-agent approach improves prediction accuracy and handles cold-start issues, it is computationally intensive and less suitable for real-time or resource-constrained environments.
@@ 4.1. Natural Language User Profile Generation
Following the approach of (Li et al., 2024) and using their prompt formatting structure, user check-in records and trajectories were first transformed into prompts. Trajectory prompting (Radford et al., 2021; Xue et al., 2022; Zhou et al., 2022) allows the conversion of heterogeneous data from LBSNs into natural language sentences that LLMs can understand. These trajectory prompts were then concatenated and used to generate NL user profiles, as outlined in Table 1. In addition to generating a 200-word NL user profile, the LLM was tasked with predicting the 3 components above. To facilitate easier parsing, the output is structured in JSON format. Examples of user profiles can be found in Appendix D.
@@ 4.2. Trajectory Prompting with Natural Language User Profiles
The NL user profile generated using the method described in §4.1 serves as the system prompt for the LLM, providing a description of the user that the model is intended to simulate. Given that the output is formatted in JSON, with each component represented as a different key, it can be easily parsed and integrated into the input prompt. This system prompt informs the LLM of the user’s attributes, including age, gender, educational background, socioeconomic level, BFI traits, preferences, and routines. Consequently, the LLM’s POI prediction output can leverage the insights provided by the NL user profile. Additionally, the LLM can learn trajectory patterns and routines from users with similar NL profiles, thereby enhancing its predictive capabilities and helping to mitigate cold-start problems.
The current trajectory block consists of all check-in records from the current session, except for the latest record, which is placed in the target block. Unlike LLM4POI, we excluded historical check-in records from the user and other users to isolate the impact of our proposed NL user profiles and minimize the input context length. The range of POI IDs specified in the instruction block helps guide the LLM in generating a predicted POI ID, ensuring it falls within the set of POIs $P$. The target block, which includes the next timestamp, user ID, and next POI ID, serves as the generation targets/labels and is excluded from the input during the prediction process.
@@ 4.3. Supervised Fine-tuning
Like LLM4POI, we applied LoRA (Hu et al., 2021) as the parameter-efficient fine-tuning (PEFT) technique used during training, which involves freezing all parameters in the pre-trained LLM and only training the newly injected low-rank decomposition matrices. These matrices are subsequently merged into the pre-trained LLM during the inference process. In their setup, LoRA matrices of rank 8 were injected only into the attention layers (query, key, value, and output projection layers) were injected with the new LoRA matrices of rank 8. This significantly reduces the GPU memory requirements when fine-tuning an LLM while maintaining similar performances to full fine-tuning.
@@ 5.1.1. Datasets
Furthermore, we recreated subsets of the Global-scale Check-in Dataset (Yang et al., 2019, 2022b), also derived from FourSquare, following the pre-processing methods outlined by (Feng et al., 2024). Trajectories were grouped into 72-hour sessions, with users having fewer than 5 sessions and sessions with fewer than 4 check-ins removed. The data was split into 70% for training, 10% for validation, and 20% for testing. Only users with between 3 and 50 sessions were included in the testing set. While they only used 200 test samples due to cost constraints, we evaluated on the entire testing set. We applied this pre-processing procedure to the cities of Moscow and São Paulo, two of the largest cities included in their evaluation. However, as shown in Appendix A, our reproduction of their dataset is not identical, as they filtered trajectories using an unspecified minimum distance from the city centers. In contrast, we used actual geometric boundaries of the cities obtained from OpenStreetMap [^5] https://www.openstreetmap.org and included only the POIs located within those boundaries. The actual geometric boundaries of cities are provided in the map located in Appendix B. All of the datasets we used are publicly available and anonymous.
@@ 5.1.2. Baselines
  • LLM-ZS (Beneduce et al., 2024): A more simplified version of LLM-Mob, also utilizes historical and contextual trajectories, simpler prompting, and zero-shot reasoning. Compares zero-shot, one-shot, and few-shot prompting predictions.
@@ 5.1.3. Evaluation
We evaluated our model using Accuracy@1, the same metric employed by both LLM4POI and AgentMove, to assess the accuracy of predicting the next immediate POI.
@@ 5.1.5. Implementation
Our fine-tuning implementation details are provided in Appendix C, based on the setup from (Li et al., 2024). We used a linearly decaying learning rate scheduler for all fine-tuning runs, with 20 warmup steps and no weight decay. Since we excluded historical trajectories from our prompts, we used a context length of 16,384 tokens and trained for 3 epochs. The Transformer Reinforcement Learning (TRL) library (von Werra et al., 2020; Wolf et al., 2020) was used throughout our SFT experiments, which were conducted on NVIDIA H100 GPUs.
@@ 5.2. Results
Table 3. Evaluation results of our proposed method incorporating natural language user profiles as system prompts for LLM-based next POI prediction on New York City, Tokyo, and California datasets.
@@
Moreover, the larger GenUP-Llama-3.1-8B model, further increases performance, achieving the highest accuracy in all cities (NYC: 0.2582, TKY: 0.2127, CA: 0.1339). Notably, this model demonstrated competitive results compared to GETNext which exploits collaborative signals from cross-user data. Similarly, the smaller GenUP-Llama-3.2-1B model also consistently surpasses the LLM4POI* baseline, showing an improvement in accuracy despite using a smaller yet more powerful pre-trained base LLM.
Table 4. Evaluation results of our method compared to in-context learning and agentic LLM frameworks for next POI prediction on Moscow and São Paulo datasets. Results marked with are based on different dataset versions of the same city.
@@ 5.3.1. Impact of Natural Language User Profile Components
Table 5. Ablation study of the different natural language user profile components as system prompts on the NYC dataset.
@@ 5.3.3. Efficiency Comparison of Trajectory Similarity and User Profiling
In contrast, our method replaces frequent trajectory similarity computations by utilizing user profiles that require only a single inference per user given their historical trajectories $\mathcal{T}_{u}$. These profiles can be easily updated periodically as new trajectories become available, ensuring they reflect evolving user behaviors without constant recalibration. Consequently, the time complexity for generating these profiles is only $O(|U|)$ where $|U|$ is the number of users and $|U|\ll L$ This significant reduction in computational overhead, coupled with the ability to adapt efficiently to changes in user preferences, positions our method as a scalable solution for recommendation systems. Additionally, omitting historical trajectories from our prompt reduces the context length to 16,384 tokens, which is half of what LLM4POI required.
@@ 5.3.4. Open-source and Lightweight Base LLM
As shown in Table 7, our approach demonstrates competitive results compared to other LLM-based methods, all while utilizing open-source and lightweight models. Remarkably, the GenUP-Llama-3.2-1B model exhibited impressive predictive capabilities despite its smaller size, making it a promising candidate for offline or on-device POI recommendations, such as on mobile devices. Importantly, the reported results of LLM-ZS, LLM-Mob, and AgentMove in Table 4 used the closed-source GPT-4o Mini as their base LLM. Originally, LLM-Mob utilized GPT-3.5 Turbo (Wang et al., 2023), whereas LLM-ZS requires at least a 70 billion, instruction-tuned LLM or closed-source models like GPT-3.5, GPT-4, or GPT-4o to start exhibiting significant advantages (Beneduce et al., 2024). Although we also employed GPT-4o mini as our NL user profile generator in the experiments, this step can easily be replaced with open-source alternatives, such as instruction-tuned variants of the Llama 3.2 models that offer similar generation capabilities.

### v2

@@ Abstract.
Traditional Point-of-Interest (POI) recommendation systems often lack transparency, interpretability, and scrutability due to their reliance on dense vector-based user embeddings. Furthermore, the cold-start problem—where systems have insufficient data for new users—limits their ability to generate accurate recommendations. Existing methods often address this by leveraging similar trajectories from other users, but this approach can be computationally expensive and increases the context length for LLM-based methods, making them difficult to scale. To address these limitations, we propose a method that generates natural language (NL) user profiles from large-scale, location-based social network (LBSN) check-ins, utilizing robust personality assessments and behavioral theories. These NL profiles capture user preferences, routines, and behaviors, improving POI prediction accuracy while offering enhanced transparency. By incorporating NL profiles as system prompts to LLMs, our approach reduces reliance on extensive historical data, while remaining flexible, easily updated, and computationally efficient. Our method is not only competitive with other LLM-based and complex agentic frameworks but is also more scalable for real-world POI recommender systems. Results demonstrate that our approach consistently outperforms baseline methods, offering a more interpretable and resource-efficient solution for POI recommendation systems. Our source code is available at: https://github.com/w11wo/GenUP/.
@@ 1. Introduction
Location-based social networks (LBSNs) generate vast amounts of data, offering significant potential for modeling human mobility patterns. One application of LBSN data is next Point-of-Interest (POI) recommendation, where the goal is to predict a user’s likely next destination based on their historical check-ins. POI recommendation systems have evolved from early Markov models (Cheng et al., 2013; Rendle et al., 2010) to more advanced deep learning techniques, such as sequential models (Kong and Wu, 2018; Sun et al., 2020) and Graph Neural Networks (Lim et al., 2020; Zhang et al., 2022; Rao et al., 2022; Yang et al., 2022a). These models have significantly improved the ability to capture spatial-temporal dependencies and user mobility patterns. However, challenges remain, particularly with the cold-start problem, where limited interaction data makes it difficult to provide accurate predictions for inactive users. This is typically resolved by learning from similar historical trajectories from other users. More recently, large language models (LLMs) have been applied to next POI recommendation (Wang et al., 2023; Beneduce et al., 2024), leveraging their commonsense reasoning capabilities and their ability to incorporate contextual information. However, LLM-based methods can be computationally intensive due to their reliance on long historical trajectory prompts to capture cross-user patterns (Li et al., 2024). Similarly, agentic LLM frameworks (Feng et al., 2024), which employ multiple agents for different tasks introduce added complexity and resource demands.
@@
Another challenge in POI recommendation is personalization, often achieved through dense vector representations of users. While these embeddings effectively model user behavior, they lack interpretability, transparency, and scrutability—qualities crucial for building user trust and improving recommendation explainability (O’Donovan and Smyth, 2005; Zhang and Chen, 2020). Natural language (NL) user profiles offer a promising alternative, where new users can quickly provide a concise description of their preferences, and existing users can update their profiles for more personalized and transparent recommendations (Radlinski et al., 2022). Despite their potential, NL profiles have not seen widespread adoption due to difficulties in evaluation and the ongoing development of methods for accurately generating these profiles (Ramos et al., 2024). Prior work, such as LLMob (Wang et al., 2024a), relies on rigid user categorization, generating multiple candidate personas and selecting the most likely one based on a given trajectory. While effective, this approach limits personalization by constraining users to predefined categories.
@@
In response, we introduce NL user profiles for next POI recommendation, replacing long historical trajectory prompts with concise profiles to reduce computational overhead and context length. We propose generating these profiles from LBSN check-ins to learn user behaviors and integrate robust behavioral theories for more accurate POI predictions. Our contributions are as follows:
@@
  • We introduce natural language (NL) user profiles as LLM system prompts for next POI recommendation.
  • We reconstruct NL user profiles from LBSN data, capturing individualized routines without predefined categories.
  • We combine LLM generation and behavioral theories for more grounded user profiles.
Our method enhances POI recommendation by improving transparency, reducing reliance on extensive historical data, and maintaining competitive accuracy with a more scalable, resource-efficient design. Unlike existing LLM-based approaches, it achieves strong performance while using smaller, open-source models, making it more adaptable for real-world deployment.
@@ 2.1. Large Language Models for Next POI Recommendation
Large language models (LLMs) have since emerged as a viable solution for next POI recommendation, offering the ability to process both spatial-temporal data and commonsense knowledge. LLM-Mob (Wang et al., 2023), for example, introduced in-context learning and exploited the zero-shot reasoning capabilities of LLMs (Wei et al., 2022) to enhance next POI recommendation given the historical and contextual trajectories of a user. Similarly, LLM-ZS (Beneduce et al., 2024) simplified the zero-shot process and analyzed zero-shot, one-shot, and few-shot prompting scenarios. However, both methods depend heavily on task-specific prompting and require very large LMs, such as 70B instruction-tuned or closed-source LLMs, to achieve competitive performance. AgentMove (Feng et al., 2024) extends these techniques by employing multiple LLMs agents (Wang et al., 2024b; Xi et al., 2023), each dedicated to different components of mobility prediction. This framework further integrates external knowledge sources such as urban structures, geospatial data, and location graphs, enabling the predictor LLM to process richer contextual information for the prediction process. While this multi-agent approach improves prediction accuracy and handles cold-start issues, it is computationally intensive and less suitable for real-time or resource-constrained environments.
@@ 4.1. Natural Language User Profile Generation
Following the approach of (Li et al., 2024) and using their prompt formatting structure, user check-in records and trajectories were first transformed into prompts. Trajectory prompting (Radford et al., 2021; Xue et al., 2022; Zhou et al., 2022) allows the conversion of heterogeneous data from LBSNs into natural language sentences that LLMs can understand. These trajectory prompts were then concatenated and used to generate NL user profiles, as outlined in Table 1. In addition to generating a 200-word NL user profile, the LLM was tasked with predicting the 3 components above. To facilitate easier parsing, the output is structured in JSON format. Examples of a user profile can be found in Appendix A.
Validation and Challenges
A key challenge in user profile generation is validation, as extensively discussed in §2.2. Another challenge we encountered was LLM hallucinations, where it failed to follow the instructed JSON schema, leading to inconsistent parsing due to changed JSON key names. However, this issue affected fewer than five users and was easily resolved by re-running the prompt.
To validate our generated profiles, we ensured that predicted personality traits were valid BFI categories and that user attributes matched predefined classes. Additionally, we employed the LLM-as-a-Judge method (Zheng et al., 2023), where LLMs assessed the quality of generated profiles. While this approach has potential biases, we consider it a reasonable proxy for validation and have provided further details in Appendix B. Interestingly, our user profiles are also weak in conciseness, parallel with the findings of (Ramos et al., 2024), who employed human evaluators. To further evaluate the biases in our generated NL user profiles, we present their distribution in Appendix C.
@@ 4.2. Trajectory Prompting with Natural Language User Profiles
The NL user profile generated using the method described in §4.1 serves as the system prompt for the LLM, providing a description of the user that the model is intended to simulate. Given that the output is formatted in JSON, with each component represented as a different key, it can be easily parsed and integrated into the input prompt. This system prompt informs the LLM of the user’s attributes, including age, gender, educational background, socioeconomic level, BFI traits, preferences, and routines. Consequently, the LLM’s POI prediction output can leverage the insights provided by the NL user profile. Additionally, the LLM can learn trajectory patterns and routines from users with similar NL profiles, thereby enhancing its predictive capabilities.
The current trajectory block consists of all check-in records from the current session, except for the latest record, which is placed in the target block. Unlike LLM4POI, we excluded historical check-in records from the user and other users to isolate the impact of our proposed NL user profiles and minimize the input context length. [^2] Our method only implicitly incorporates historical information into its user profile, unlike LLM4POI. The latter has access to multiple similar, repeatable trajectories, whereas the former can only summarize frequent or long-term trajectories into a concise user profile. LLM4POI also incorporates intra-user social data, while ours does not. The range of POI IDs specified in the instruction block helps guide the LLM in generating a predicted POI ID, ensuring it falls within the set of POIs $P$. The target block, which includes the next timestamp, user ID, and next POI ID, serves as the generation targets/labels and is excluded from the input during the prediction process.
@@ 4.3. Supervised Fine-tuning
Like LLM4POI, we applied LoRA (Hu et al., 2021) as the parameter-efficient fine-tuning (PEFT) technique used during training, which involves freezing all parameters in the pre-trained LLM and only training the newly injected low-rank decomposition matrices. These matrices are subsequently merged into the pre-trained LLM during the inference process. In their setup, LoRA matrices of rank 8 were injected into the attention layers (query, key, value, and output projection layers). This significantly reduces the GPU memory requirements when fine-tuning an LLM while maintaining similar performances to full fine-tuning.
@@ 5.1.1. Datasets
Furthermore, we recreated subsets of the Global-scale Check-in Dataset (Yang et al., 2019, 2022b), also derived from FourSquare, following the pre-processing methods outlined by (Feng et al., 2024). Trajectories were grouped into 72-hour sessions, with users having fewer than 5 sessions and sessions with fewer than 4 check-ins removed. The data was split into 70% for training, 10% for validation, and 20% for testing. Only users with between 3 and 50 sessions were included in the testing set. While they only used 200 test samples due to cost constraints, we evaluated on the entire testing set. We applied this pre-processing procedure to the cities of Moscow and São Paulo, two of the largest cities included in their evaluation. However, as shown in Appendix D, our reproduction of their dataset is not identical, as they filtered trajectories using an unspecified minimum distance from the city centers. In contrast, we used actual geometric boundaries of the cities obtained from OpenStreetMap [^6] https://www.openstreetmap.org and included only the POIs located within those boundaries. All of the datasets we used are publicly available and anonymous.
@@ 5.1.2. Baselines
  • LLM-ZS (Beneduce et al., 2024): A simplified version of LLM-Mob, also utilizes historical and contextual trajectories, simpler prompts, and zero-shot reasoning. Compares zero-shot, one-shot, and few-shot prompting predictions.
@@ 5.1.3. Evaluation
We evaluated our model using Accuracy@1, the same metric employed by both LLM4POI and AgentMove, to assess prediction accuracy when a user can only visit one next POI rather than multiple locations.
@@ 5.1.5. Implementation
Our fine-tuning implementation details are provided in Appendix E, based on the setup from (Li et al., 2024). We used a linearly decaying learning rate scheduler for all fine-tuning runs, with 20 warmup steps and no weight decay. Since we excluded historical trajectories from our prompts, we used a context length of 16,384 tokens and trained for 3 epochs. The Transformer Reinforcement Learning (TRL) library (von Werra et al., 2020; Wolf et al., 2020) was used throughout our SFT experiments, which were conducted on NVIDIA H100 GPUs.
@@ 5.2. Results
Table 3. Evaluation results of our proposed method incorporating natural language user profiles as system prompts for LLM-based next POI prediction on New York City, Tokyo, and California datasets. [^8] LM4POI* refers to the baseline method without historical trajectories (§5.1.4). For clarity, the highest score on each dataset is highlighted in bold, while the second highest is underlined.
@@
Moreover, the larger GenUP-Llama-3.1-8B model, further increases performance, achieving the highest accuracy in all three cities (NYC: 0.2582, TKY: 0.2127, CA: 0.1339). Notably, this model demonstrated competitive results compared to GETNext which exploits collaborative signals from cross-user data. Similarly, the smaller GenUP-Llama-3.2-1B model also consistently surpasses the LLM4POI* baseline, showing an improvement in accuracy despite using a smaller yet more powerful pre-trained base LLM.
Table 4. Evaluation results of our method compared to in-context learning and agentic LLM frameworks for next POI prediction on Moscow and São Paulo datasets. Results marked with are based on different dataset versions of the same city.footnotemark:
@@ 5.3.1. Impact of Natural Language User Profile Components
Table 5. Ablation study of the different natural language user profile components as system prompts on the NYC dataset.footnotemark:
@@ 5.3.2. User Cold-start Analysis
We also analyzed how different NL user profile components affect different user groups, with results shown in Table 7. Our findings indicate that their impact varies by user activity level. For inactive (cold-start) users, incorporating routines and preferences improved accuracy by +0.93%, with user attributes providing another +0.93% boost. However, while adding BFI traits reduced accuracy for inactive users – as also observed in §5.3.1 – they enhanced the performance for highly active users.
@@
Table 7. Impact of different NL user profile components on next POI recommendation accuracy for NYC users, categorized by activity level, using the GenUP-Llama2-7B model.
System Prompt Inactive Normal Very Active
200-word User Profile 0.2012 0.2737 0.2728
+ Routines & Preferences 0.2105 0.2774 0.2680
+ User Attributes 0.2198 0.2664 0.2692
+ BFI Traits 0.2105 0.2591 0.2752
@@ 5.3.3. Efficiency Comparison of Trajectory Similarity and User Profiling
In contrast, our method replaces frequent trajectory similarity computations by utilizing user profiles that require only a single inference per user given their historical trajectories $\mathcal{T}_{u}$. Consequently, the time complexity for generating these profiles is only $O(|U|)$ where $|U|$ is the number of users and $|U|\ll L$. These profiles can be easily updated periodically as new trajectories become available, ensuring they reflect evolving user behaviors without constant recalibration. An example is to employ a user profile embedding shift, where a new profile $P^{\prime}$ replaces an existing profile $P$ if their similarity falls below a predefined threshold, indicating a significant behavioral shift. This strategy prioritizes meaningful long-term changes over fixed update intervals while allowing short-term adaptations when necessary. In practice, user profile generation with GPT-4o Mini averaged under 2 seconds per user and can be further accelerated with batched and parallel inference, making updates efficient even at scale. Notably, our method omits historical trajectories halving the context length of LLM4POI. This significantly reduces memory and compute requirements.
@@ 5.3.4. Open-source and Lightweight Base LLM
As shown in Appendix F, our approach demonstrates competitive results compared to other LLM-based methods, all while utilizing open-source and lightweight models. Remarkably, the GenUP-Llama-3.2-1B model exhibited impressive predictive capabilities despite its smaller size, making it a promising candidate for offline or on-device POI recommendations, such as on mobile devices. Importantly, the reported results of LLM-ZS, LLM-Mob, and AgentMove in Table 4 used the closed-source GPT-4o Mini as their base LLM. Originally, LLM-Mob utilized GPT-3.5 Turbo (Wang et al., 2023), whereas LLM-ZS requires at least a 70 billion, instruction-tuned LLM or closed-source models like GPT-3.5, GPT-4, or GPT-4o to start exhibiting significant advantages (Beneduce et al., 2024). Although we also employed GPT-4o mini as our NL user profile generator in the experiments, this step can easily be replaced with open-source alternatives, such as instruction-tuned variants of the Llama 3.2 models that offer similar generation capabilities.
6. Conclusion and Future Work
Acknowledgements.
We would like to thank the support of the National Computational Infrastructure and the ARC Center of Excellence for Automated Decision Making and Society (CE200100005). Computational facilities were provided by the School of Computer Science and Engineering at UNSW Sydney through the Wolfpack computational cluster. We further thank OpenAI’s Researcher Access Program for API access to GPT models and the Google Cloud Research Credits program for their support.
Appendix A Sample User Profile Generation Results
We provide a sample generation result of a user profile from the NYC dataset in Figure 2. We used GPT-4o Mini to generate profiles given the check-in records of each user.
Appendix B User Profile Validation
We followed (Ramos et al., 2024) and their qualitative case study on NL profiles, adopting their metrics to assess our generated user profiles. LLMs were prompted using the validation prompt shown in Table 8, with evaluation results presented in Table 9.
Table 8. User profile validation prompt provided to an LLM for evaluation, using metrics from (Ramos et al., 2024).
prompt Given the following natural language (NL) user profile [user profile], you are to assess the user profile based on four criteria:
- Fluency: Is the NL profile both syntactically and semantically correct?
- Informativeness: Does the NL profile provide important information for a user profile?
- Conciseness: Is the NL profile written in a concise manner?
- Relevance: Given the list of reviews, is the NL profile relevant to the user?
Return your response in the JSON format: {”fluency”: 0/1, ”informativeness”: 0/1, ”conciseness”: 0/1, ”relevance”: 0/1}
Table 9. Evaluation results of LLM-judged user profile validation using metrics from (Ramos et al., 2024).
Metric NYC CA TKY
Judge: GPT 4-o Mini
Fluency 100.00 100.00 100.00
Informativeness 100.00 100.00 100.00
Conciseness 43.17 37.26 40.60
Relevance 100.00 100.00 100.00
Judge: Gemini 1.5 Flash
Fluency 100.00 100.00 100.00
Informativeness 100.00 100.00 100.00
Conciseness 65.81 57.68 63.48
Relevance 100.00 100.00 100.00
Appendix C User Profile Distribution
We present the distribution of the generated NL user profiles across three datasets: NYC, CA, and TKY. Table 10 displays the distribution of predicted personality traits, while Table 11 shows the distribution of user attributes, highlighting the diversity within each dataset.
Table 10. Distribution of predicted BFI traits across the NYC, CA, and TKY datasets.
NYC CA TKY
Extroverted 736 3526 1476
Introverted 311 430 805
Agreeable 1047 3956 2281
Antagonistic 0 0 0
Conscientious 1047 3956 2281
Unconscientious 0 0 0
Neurotic 41 268 99
Emotionally Stable 1006 3688 2182
Open to Experience 1043 3947 2279
Closed to Experience 4 9 2
Table 11. Distribution of predicted user attributes across the NYC, CA, and TKY datasets.
NYC CA TKY
Age Child 1 14 2
Adolescent 72 330 274
Adult 974 3612 2005
Older Adult 0 0 0
Gender Male 774 2911 1881
Female 273 1045 400
Education Some Schooling 43 190 192
High School 55 176 165
College & Beyond 949 3590 1924
Socioeconomic Lower 3 7 3
Middle 1028 3651 2240
Upper 16 298 38
Appendix D Dataset Statistics
City Users Trajectories POIs Check-ins
AgentMove (Feng et al., 2024)
Moscow 10,501 100,854 93,599 950,898
São Paulo 11,856 77,120 78,904 809,198
Ours
Moscow 6,864 66,216 89,247 500,213
São Paulo 3,718 40,045 64,690 300,044
Appendix E Hyperparameters
Model Learning Batch Size
Rate NYC Others
GenUP-Llama-2-7B 2e-5 1 2
GenUP-Llama-3.1-8B 2e-4 1 2
GenUP-Llama-3.2-1B 2e-4 1 2
Table 14. LoRA parameters based on LLM4POI (Li et al., 2024).
parameter value
rank 8
alpha 16
dropout 0.0
bias none
modules query, key, value, output
Appendix F Computational Efficiency Comparison
In Table 15, we compare the base LLMs used in our approach with those employed by existing methods. Our method leverages open-source, lightweight models, while previous works such as LLM-ZS (Beneduce et al., 2024), LLM-Mob (Wang et al., 2023), and AgentMove (Feng et al., 2024) rely on larger and/or or closed-source LLMs.